OpenStudy (anonymous):

Please help me!! This are piecewise functions. Evaluate the following for the given function. f(x)={ 3x-7, if x is less than or equal to 2 {6-2x, if x>2 1. f(2) 2. f(-3) 3. f(5)

OpenStudy (anonymous):

f(x)= [3x-7, x \le2\] \[6-2x, x >2\] You would have to plug in the f(#) into the equation that it falls into. So for f(2), you would use the first equation since x is less than or equal to 2. 3(2)-7 =6-7 =-1

OpenStudy (anonymous):

Ok, I get it now. Thank you!

OpenStudy (anonymous):

You're welcome:)